OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y greater than or equal to -4x - 2? (-1, 0) (0, 0) (4, 4) (1, 3)

OpenStudy (anonymous):

When you do this type of problem, just remember that the first number is the x-coordinate, and the second number is the y-coordinate. For each number, you're going to see if the inequality is true or false when you make a substitution. Let's take (-1, 0) as an example. \[y \ge -4x-2\]\[0 \ge -4(-1)-2\]\[0 \ge 4-2\]\[0 \ge 2\]but zero isn't larger than 2, so this is false. This means that (-1, 0) doesn't work. You have to go through and check each of those 4. There could be multiple right answers.

OpenStudy (anonymous):

so are you saying (1,0) is the correct answer ?

OpenStudy (anonymous):

I'm saying that (-1, 0) is not correct.

OpenStudy (anonymous):

ok. can you help me with another one please?

OpenStudy (anonymous):

@jabberwock

OpenStudy (anonymous):

Sure thing.

OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y greater than or equal to 4x - 2? (0, -4) (0, 0) (0, -3) (0, -5)

OpenStudy (anonymous):

You're going to do this one the same way. Did you finish the last one? (You have to check all 4 because there might be multiple correct answers)

OpenStudy (anonymous):

i dont understand how to check them

OpenStudy (anonymous):

Let's look back at the example of (-1, 0). -1 is an x-coordinate, and 0 is a y-coordinate, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. So now you're going to take your x-coordinate and your y-coordinate, and you're going to plug them into your inequality. \[y \ge -4x-2\]\[0 \ge -4(-1) -2\] Okay with that part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, so if we do the arithmetic and simplify it all down, we get \[0 \ge 2\]is that a true statement or a false statement?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Right. Now, since it's false, we know it doesn't work. If we would have gotten a true statement (something like 17>=0), then we would have said "yes." Make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okidoki. Sorry to do this, but I have to get going. Maybe try reposting one of the questions if you get stuck. Good luck!